# The Double-Ellipsoid Geometry of CLIP

### v2

@@ Abstract
Contrastive Language-Image Pre-Training (CLIP) is highly instrumental in machine learning applications within a large variety of domains. We investigate the geometry of this embedding, which is still not well understood, and show that text and image reside on linearly separable ellipsoid shells, not centered at the origin. We explain the benefits of having this structure, allowing to better embed instances according to their uncertainty during contrastive training. Frequent concepts in the dataset yield more false negatives, inducing greater uncertainty. A new notion of conformity is introduced, which measures the average cosine similarity of an instance to any other instance within a representative data set. We show this measure can be accurately estimated by simply computing the cosine similarity to the modality mean vector. Furthermore, we find that CLIP’s modality gap optimizes the matching of the conformity distributions of image and text.
@@ 1 Introduction
Leveraging this deeper understanding, we introduce a new definition of concept conformity, quantifying how close a sample resides with respect to all others. We find a remarkable correlation between conformity and cosine similarity to the mean vector (Pearson correlation: 0.9998 for MS-COCO). Furthermore, we show that the distribution of conformity differs between modalities, with CLIP’s ellipsoid alignment offering a plausible explanation for the modality gap.
@@ 6.1 Conformity
In Fig. 9, $C$ versus $\hat{C}$ are plotted for the entire MS-COCO set, for both image and text embeddings. A close to perfect correlation is obtained, with Pearson correlation of 0.9998 for both image and text where $a=1.411$, $b=-0.008$ for text and $a=1.461$, $b=-0.002$ for images.
@@ Appendix B Statistical Analysis
Appendix C Additional Experiments and Visualizations
C.1 Conformity
C.2 Reaffirming loss and conformity matching experiments

### v3

@@ Abstract
Contrastive Language-Image Pre-Training (CLIP) is highly instrumental in machine learning applications within a large variety of domains. We investigate the geometry of this embedding, which is still not well understood, and show that text and image reside on linearly separable ellipsoid shells, not centered at the origin. We explain the benefits of having this structure, allowing to better embed instances according to their uncertainty during contrastive training. Frequent concepts in the dataset yield more false negatives, inducing greater uncertainty. A new notion of conformity is introduced, which measures the average cosine similarity of an instance to any other instance within a representative data set. We prove this measure can be accurately estimated by simply computing the cosine similarity to the modality mean vector. Furthermore, we find that CLIP’s modality gap optimizes the matching of the conformity distributions of image and text.
@@ 1 Introduction
Leveraging this deeper understanding, we introduce a new definition of concept conformity, quantifying how close a sample resides with respect to all others. We prove that there is a proportion between conformity and cosine similarity to the mean vector (See proof in Supp. C1, and empirically with Pearson correlation: 0.9998 for MS-COCO). Furthermore, we show that the distribution of conformity differs between modalities, with CLIP’s ellipsoid alignment offering a plausible explanation for the modality gap.
@@ 6.1 Conformity
In Appendix C1 we prove this correlation under the thin-shell assumption, and in Fig. 9, $C$ versus $\hat{C}$ are plotted for the entire MS-COCO set, for both image and text embeddings. A close to perfect correlation is obtained, with Pearson correlation of 0.9998 for both image and text where $a=1.411$, $b=-0.008$ for text and $a=1.461$, $b=-0.002$ for images, validating with close to perfect alignment with the rigorous mathematic derivation.
@@ Appendix B Statistical Analysis
Appendix C Additional Experiments and Visualizations
C.1 Close relations between conformity and surrogate-conformity
We show below the validity of our conformity approximation under the thin-shell assumption.
Proposition 1.
Let $S=\{v^{1},\dots,v^{N}\}$ be a set of $N$ vectors in $\mathbb{R}^{F}$ exhibiting the thin-shell phenomenon, i.e.,
$\|v^{i}-\bar{v}\|\approx R\quad\text{for all }i,$
where $\bar{v}=\frac{1}{N}\sum_{j=1}^{N}v^{j}$ is the sample mean and we use the Euclidean norm $\|v^{i}\|^{2}=\sum_{k}(v^{i}_{k})^{2}$. Then, for any $v^{j}\in S$, the following approximation holds:
$\mathbb{E}_{v^{j}\in S}[\cos(v^{i},v^{j})]\approx A\cdot\cos(v^{i},\bar{v}),$ (13)
where $A\approx\frac{\sqrt{\mu_{\text{norm}}^{2}+R^{2}}}{\mu_{\text{norm}}}$, $\mu_{\text{norm}}=\|\bar{v}\|$ and the symbol $\approx$ represents the shell approximation (which becomes more accurate as the width of the shell decreases) and approximate orthogonality between a random vector and the mean vector.
Proof.
We start by expanding the left-hand side:
$\mathbb{E}_{v^{j}\in S}[\cos(v^{i},v^{j})]=\frac{1}{N}\sum_{j=1}^{N}\frac{v^{i}\cdot v^{j}}{\|v^{i}\|\cdot\|v^{j}\|}.$
Writing explicitly the inner-product we have:
$\frac{1}{N\cdot\|v^{i}\|}\sum_{j=1}^{N}\frac{1}{\|v^{j}\|}\sum_{k=1}^{F}v^{i}_{k}v^{j}_{k}.$
Now, consider the right-hand side of Equation (13):
$\cos(v^{i},\bar{v})=\frac{v^{i}\cdot\bar{v}}{\|v^{i}\|\cdot\|\bar{v}\|}=\frac{1}{\|v^{i}\|\cdot\mu_{\text{norm}}}\sum_{k=1}^{F}v^{i}_{k}\left(\frac{1}{N}\sum_{j=1}^{N}v^{j}_{k}\right)=\frac{1}{N\cdot\|v^{i}\|\cdot\mu_{\text{norm}}}\sum_{j=1}^{N}\sum_{k=1}^{F}v^{i}_{k}v^{j}_{k}.$
Observe that the only difference between the two expressions lies in the difference between $\mu_{\text{norm}}$ and $\|v^{j}\|$. We show below that under the thin-shell assumption $\|v^{j}\|\approx\sqrt{R^{2}+\mu_{\text{norm}}^{2}}$.
Let us define by $z^{j}$ the difference vector between a vector $v^{j}$ and the mean vector $\bar{v}$, that is $z^{j}=v^{j}-\bar{v}$. Then,
$\|v^{j}\|^{2}=\|z^{j}+\bar{v}\|^{2}=\|z^{j}\|^{2}+2z^{j}\cdot\bar{v}+\|\bar{v}\|^{2}.$
In high dimensions, the inner product $z^{j}\cdot\bar{v}$ is small due to approximate orthogonality, so:
$\|v^{j}\|^{2}\approx\|z^{j}\|^{2}+\mu_{\text{norm}}^{2}\approx R^{2}+\mu_{\text{norm}}^{2}.$
Taking square roots:
$\|v^{j}\|\approx\sqrt{R^{2}+\mu_{\text{norm}}^{2}}.$
Thus, the scalar factor $A$ in Equation (13) is given by:
$A=\frac{\mu_{\text{norm}}}{\|v^{j}\|}\approx\frac{\mu_{\text{norm}}}{\sqrt{R^{2}+\mu_{\text{norm}}^{2}}}.$
Empirically we know for Vit-B/32 that $\mu_{norm}=7.587$ and $R\approx 7.59$, thus the mathematical derivation state that $A^{-1}=\frac{\sqrt{7.59^{2}+7.587^{2}}}{7.587}=1.414$ For images and $A^{-1}=\frac{\sqrt{5.59^{2}+5.75^{2}}}{5.75}=1.4$, very close to the empirical observations (note that the correlation is reversed in the main paper).
C.2 Conformity
@@
C.3 Reaffirming loss and conformity matching experiments